40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 6 19:53:15 2013, Jenne, Update, LSC, Arms locked in IR, aligned. IFO at nominal power 
    Reply  Wed Aug 7 10:09:30 2013, Koji, Update, LSC, Arms locked in IR, aligned. IFO at nominal power 
       Reply  Wed Aug 7 21:52:11 2013, Jenne, Update, SUS, SRM coils fine - problem with slow bias slider 
          Reply  Fri Aug 9 21:05:28 2013, Koji, Update, SUS, fixed: SRM coils fine - problem with slow bias slider 
Message ID: 8975     Entry time: Wed Aug 7 10:09:30 2013     In reply to: 8974     Reply to this: 8981
Author: Koji 
Type: Update 
Category: LSC 
Subject: Arms locked in IR, aligned. IFO at nominal power 

I have a concern about the SRM suspension. The yaw alignment bias produces huge pitch coupling.

This could be a connector issue or the rubbing of the mirror on the EQ stops.

We have the photos of the magnets and they were not touching the OSEMs.

ELOG V3.1.3-